### v4

# Hamiltonian Active Particles in Incompressible Fluid Membranes

Abstract
Active proteins and membrane-bound motors exert force dipole flows along fluid interfaces and lipid bilayers. We develop a Hamiltonian framework for the interactions of pusher and puller dipoles embedded in an incompressible two-dimensional membrane supported by a shallow viscous subphase. Beginning from the Brinkman-regularized Stokes equations of the membrane–subphase system, we construct the near- and far-field dipolar velocity and associated stream functions. Although generic dipoles reorient under the local membrane vorticity, we show that the far-field dipolar flow is vorticity-free; force-free motors therefore retain fixed orientation and obey a position-based Hamiltonian dynamics in which the positions of $N$ dipoles evolve via an effective Hamiltonian built from the dipolar stream function. In the near field, where the flow possesses finite vorticity, a Hamiltonian formulation is recovered in the quenched-orientation limit. We demonstrate this Hamiltonian structure by simulating ensembles of pusher and puller dipoles and compare the dynamics generated by the near-field and far-field Hamiltonians. For identical dipoles, the far-field Hamiltonian produces rapid clustering from random initial conditions, whereas the near-field Hamiltonian suppresses collapse and yields extended, non-aggregating configurations. Our work thus provides a concrete realization of position-based Hamiltonian descriptions for active particles in incompressible fluid membranes and establishes the important roles of membrane incompressibility and fluid vorticity that leads to the emergence of such Hamiltonian descriptions.
I Introduction
Active proteins, enzymes, and membrane–bound motors continually inject stress into biological interfaces, driving flows along fluid membranes and lipid bilayers. At the micron scale, inertia is negligible and these inclusions generate Stokesian flow fields whose leading contribution is a force dipole (stresslet) rather than a monopole [1, 2]. Such dipoles—extensile “pushers” and contractile “pullers”—play central roles in pattern formation, instabilities, and collective behaviour across active fluids [3, 4, 5]. Membrane–bound proteins and enzymatic assemblies act as force dipoles in vivo [6, 7, 8], and experiments have demonstrated motor–mediated clustering, cooperative pulling, and tension generation in reconstituted membranes [9, 10]. When these dipoles are confined to a two–dimensional interface, their hydrodynamic couplings differ sharply from those in bulk fluids, owing to the effectively planar geometry and momentum leakage into the surrounding three–dimensional subphase.
Supported membranes provide a particularly interesting setting in which these hydrodynamic effects have been worked out in detail. Classical analyses of membrane drag by Saffman, Delbrück, and their successors [11, 12, 13, 14, 15, 16, 17] showed that flows in a viscous membrane coupled to a shallow subphase obey a Brinkman–regularized Stokes equation. The resulting Green’s function exhibits two qualitatively distinct regimes: (i) a logarithmic “near field” with unscreened in–plane hydrodynamics producing $1/r$ dipolar flows, and (ii) a “far field” in which substrate friction produces $1/r^{3}$ dipolar flows. Recent studies of swimmers and active inclusions in thin films and membranes [18, 19, 20, 21] highlight the richness of these interactions, but a concrete realization of a Hamiltonian framework for pusher and puller motors in such incompressible membranes and importantly, the role of membrane fluid vorticity, remains to be explored and clarified. The main objective of this work is to explicitly demonstrate how such a position-based Hamiltonian description naturally emerges for active particles in incompressible fluid membranes.
In this work we develop a Hamiltonian description of pusher and puller interactions in an incompressible membrane coupled to a shallow subphase. Beginning from the screened Stokes equations of the membrane–subphase system, we obtain the real–space Green’s tensor, derive its near– and far–field asymptotics, and construct the associated dipolar velocity and stream functions. Although a generic dipole reorients in response to the local membrane vorticity, we show that the far–field dipolar flow is vorticity free. Force–free motors therefore maintain fixed orientation and admit a reduced position-based Hamiltonian description. However, in the near field, where vorticity is finite, a Hamiltonian structure is recovered in the quenched–orientation limit, closely related to recent two–dimensional Hamiltonian formulations of active particles [22].
Exploiting the Hamiltonian framework, we compare the collective behaviour generated by the near-field and far-field Hamiltonians for ensembles of identical pushers and pullers. The contrast is striking: the far–field Hamiltonian produces rapid collapse into compact clusters, whereas the near–field Hamiltonian suppresses aggregation and yields extended, non–collapsing configurations. These trends persist across a wide range of initial geometries and motor types, demonstrating that the structure of the hydrodynamic kernel—not the sign of the dipole alone—determines whether a membrane–bound active ensemble aggregates or disperses.
Taken together, our results show that even in the simplest incompressible limit, the hydrodynamic screening intrinsic to supported membranes imposes strong, kernel–dependent constraints on collective organisation. This work provides a foundation for similar Hamiltonian approaches to active matter on interfaces and sets the stage for extensions to chiral membranes, curved geometries, and biologically driven active films [20, 21, 23, 24].
II Incompressible supported membrane
We consider a two–dimensional viscous membrane of shear viscosity $\eta_{s}$ lying in the plane $z=h$, supported by a Newtonian shallow subphase of viscosity $\eta$ and thickness $h$ above a rigid wall (Fig. 1). Throughout the main text we restrict to the incompressible limit: the in–plane velocity field $\mathbf{v}(\mathbf{r})$ satisfies $\nabla\!\cdot\!\mathbf{v}=0$, there is no dilatational viscosity.
Eliminating the three–dimensional subphase flow under a lubrication approximation yields an effective Brinkman friction acting on the membrane [16, 23], so that the steady Stokes balance reads
$\eta_{s}\nabla^{2}\mathbf{v}-\nabla p-\zeta_{\parallel}\mathbf{v}+\mathbf{F}=\mathbf{0},\qquad\nabla\!\cdot\!\mathbf{v}=0,$ (1)
where $p$ is the membrane pressure, $\mathbf{F}(\mathbf{r})$ is any in–plane body force density, and $\zeta_{\parallel}$ is a phenomenological friction coefficient that encodes momentum leakage into the subphase. For a shallow film one typically expects
$\zeta_{\parallel}\sim\frac{\eta}{h},$ (2)
so that the hydrodynamic screening length
$\kappa^{-1}=\sqrt{\frac{\eta_{s}}{\zeta_{\parallel}}}=\sqrt{\frac{\eta_{s}h}{\eta}}$ (3)
sets the crossover between “membrane–dominated” and “subphase–dominated” flow.
Because Eq. (1) is linear and translationally invariant in the membrane plane, the velocity response to a localized in–plane force $\mathbf{F}(\mathbf{r})=\mathbf{f}\,\delta(\mathbf{r})$ can be written in terms of a Green’s tensor $G_{ij}(\mathbf{r})$,
$v_{i}(\mathbf{r})=G_{ij}(\mathbf{r})f_{j}.$ (4)
In Fourier space, membrane incompressibility can be utilized to obtain
$\tilde{G}_{ij}(\mathbf{q})=\frac{1}{\eta_{s}(q^{2}+\kappa^{2})}\big(\delta_{ij}-\hat{q}_{i}\hat{q}_{j}\big),$ (5)
where $\hat{\mathbf{q}}=\mathbf{q}/q$. The real–space Green’s tensor is obtained from the inverse transform
$G_{ij}(\mathbf{r})=\int\!\frac{d^{2}q}{(2\pi)^{2}}\,e^{i\mathbf{q}\cdot\mathbf{r}}\,\tilde{G}_{ij}(\mathbf{q}).$ (6)
Rotational invariance implies the decomposition
$G_{ij}(\mathbf{r})=\frac{1}{2\pi\eta_{s}}\Big[A(r)\,\delta_{ij}+B(r)\,\hat{r}_{i}\hat{r}_{j}\Big],$ (7)
with $r=|\mathbf{r}|$ and $\hat{\mathbf{r}}=\mathbf{r}/r$. Explicit integration yields
$A(r)$ $=K_{0}(\kappa r)+\frac{K_{1}(\kappa r)}{\kappa r}-\frac{1}{\kappa^{2}r^{2}},$ (8)
$B(r)$ $=-K_{0}(\kappa r)-\frac{2K_{1}(\kappa r)}{\kappa r}+\frac{2}{\kappa^{2}r^{2}},$ (9)
where $K_{n}$ are modified Bessel functions of the second kind. Equations (7)–(9) constitute the two–dimensional Stokeslet for an incompressible supported membrane [16, 17].
Figure 1: Schematic of active force–dipole motors confined to a supported fluid membrane. Orange ellipses denote motors with orientations (double arrows). The light blue layer is an incompressible two–dimensional membrane of shear viscosity $\eta_{s}$, supported by a viscous subphase of viscosity $\eta$ and thickness $h$ above a rigid wall.
III Dipolar flows: near and far zones
Force–free active inclusions such as membrane–anchored motors are naturally modelled as force dipoles (stresslets). A point stresslet of strength $\sigma$ and orientation $\hat{\mathbf{d}}$ located at the origin can be represented as the derivative of the Stokeslet along the dipole axis,
$v_{i}(\mathbf{r})=\sigma\,\hat{d}_{k}\,\partial_{k}G_{ij}(\mathbf{r})\,\hat{d}_{j},$ (10)
with $\hat{\mathbf{d}}=(\cos\alpha,\sin\alpha)$. Introducing polar coordinates $(r,\theta)$ for $\mathbf{r}$, we can express the velocity in terms of radial and azimuthal components $v_{r}$ and $v_{\theta}$.
III.1 Near field: unscreened hydrodynamics
For separations much smaller than the screening length, $\kappa r\ll 1$, the Bessel functions admit the expansions
$A(r)$ $\simeq\frac{1}{4}\Big[-1-2\gamma-2\ln\!\Big(\frac{\kappa r}{2}\Big)\Big],$ (11)
$B(r)$ $\simeq\frac{1}{2},$
where $\gamma$ is Euler’s constant. Substituting these into Eq. (10) and differentiating yields a purely radial near–field flow for a dipole placed at the origin, with the oriantation angle $\alpha$,
$\mathbf{v}_{\rm near}(\mathbf{r})=\frac{\sigma}{4\pi\eta_{s}r}\,\cos\!\big[2(\alpha-\theta)\big]\,\hat{\mathbf{r}},\qquad(\kappa r\ll 1),$ (12)
characteristic of a two–dimensional stresslet with logarithmic Stokeslet flow.
An incompressible two–dimensional flow may be written in terms of a scalar stream function $\Psi(r,\theta)$ via
$v_{r}=\frac{1}{r}\,\partial_{\theta}\Psi,\qquad v_{\theta}=-\partial_{r}\Psi.$ (13)
For Eq. (12) one finds
$\Psi_{\rm near}(r,\theta)=-\frac{\sigma}{8\pi\eta_{s}}\sin\!\big[2(\alpha-\theta)\big],$ (14)
and the scalar vorticity,
$\omega_{z}^{\rm(near)}(r,\theta)=-\frac{\sigma}{2\pi\eta_{s}r^{2}}\sin\!\big[2(\alpha-\theta)\big],$ (15)
is nonzero except along the principal axes, indicating that the near–field flow is locally vortical.
III.2 Far field: screened hydrodynamics
At distances much larger than the screening length, $\kappa r\gg 1$, the exponential decay of $K_{n}(\kappa r)$ leaves only the algebraic terms in Eqs. (8)–(9),
$A(r)\simeq-\frac{1}{\kappa^{2}r^{2}},\qquad B(r)\simeq\frac{2}{\kappa^{2}r^{2}}.$ (16)
Substituting into Eq. (10) gives the far–field velocity
$\mathbf{v}_{\rm far}(\mathbf{r})=\frac{\sigma}{\pi\zeta_{\parallel}r^{3}}\Big[\cos\!\big(2\Delta\big)\,\hat{\mathbf{r}}-\sin\!\big(2\Delta\big)\,\hat{\bm{\theta}}\Big],\qquad(\kappa r\gg 1),$ (17)
where $\Delta=\alpha-\theta$ and we have used $\kappa^{2}=\zeta_{\parallel}/\eta_{s}$. The flow now decays as $r^{-3}$ and has both radial and azimuthal components.
Using the stream–function representation, one finds
$\Psi_{\rm far}(r,\theta)=-\frac{\sigma}{2\pi\zeta_{\parallel}}\frac{\sin\!\big[2(\alpha-\theta)\big]}{r^{2}},$ (18)
which matches the near–field expression at $r\sim\kappa^{-1}$. In contrast to the near zone, the far–field flow is irrotational away from the dipole singularity:
$\omega_{z}^{\rm(far)}(r,\theta)=0\quad(r>0).$ (19)
IV Hamiltonian formulation for dipoles
The incompressible membrane enjoys a useful structural simplification: for force dipoles with fixed orientations, the positional dynamics can be written in Hamiltonian form. Let the position of dipole $i$ be $\mathbf{r}_{i}=(x_{i},y_{i})$, with strength $\sigma_{i}=\pm 1$ (pusher or puller) and fixed orientation angle $\alpha_{i}$. The velocity of dipole $i$ is obtained by evaluating the dipolar flow generated by all other dipoles $j\neq i$ at $\mathbf{r}_{i}$:
$\dot{\mathbf{r}}_{i}=\sum_{j\neq i}\mathbf{v}_{j}(\mathbf{r}_{i}-\mathbf{r}_{j}).$ (20)
Using the stream function representation, this can be recast as
$\sigma_{i}\dot{\mathbf{r}}_{i}=\hat{\mathbf{z}}\times\nabla_{i}H,\qquad H=\sum_{i\neq j}\sigma_{i}\,\Psi_{j}(\mathbf{r}_{ij}),$ (21)
where $\mathbf{r}_{ij}=\mathbf{r}_{i}-\mathbf{r}_{j}$ and $\Psi_{j}$ is the stream function generated by dipole $j$ (near or far, depending on the regime of interest). Equation (21) shows that $(x_{i},y_{i})$ play the role of canonical variables with an effective Hamiltonian determined by the hydrodynamic kernel.
Figure 2: Dynamics of $N=15$ pusher (top row) and puller (bottom row) dipoles in an incompressible supported membrane. Open circles mark initial positions; filled black dots mark final positions. (a–b,e–f) Evolution under the screened far-field Hamiltonian $H_{\rm far}$: particle trajectories (a,e) and mean pair separation $\langle d_{ij}\rangle(t)$ (b,f). Both pushers and pullers exhibit strong mutual attraction and collapse into compact clusters. (c–d,g–h) Evolution under the unscreened near-field Hamiltonian $H_{\rm near}$ in the quenched-orientation limit: trajectories (c,g) and $\langle d_{ij}\rangle(t)$ (d,h). In both cases the ensembles expand and $\langle d_{ij}\rangle(t)$ grows monotonically, demonstrating that near-field interactions suppress the far-field aggregation instability.
Figure 3: Representative far–zone dynamics of a randomly initialized cluster of twelve pushers in the incompressible membrane with random initial locations (within a disc) and random orientations. Green disks indicate initial positions, red disks indicate final positions, and the grey curves trace the particle trajectories. A soft harmonic repulsion is included to regularize close encounters.
IV.1 Near–field Hamiltonian
As noted in Eq. (15), the unscreened near–field dipolar flow carries finite vorticity, so freely rotating dipoles undergo hydrodynamic reorientation. These orientational dynamics preclude a Hamiltonian written solely in terms of particle positions. A reduced Hamiltonian description is nevertheless recovered in the quenched–orientation limit, where dipole axes are externally constrained or there is a separation in orientation relaxation timescales compared to translational motion. This regime is natural for anchored “shakers” and for strongly aligned active suspensions [22]. With orientations fixed, the near–field stresslet retains its $1/r$ form and generates a consistent position-based Hamiltonian, enabling a direct comparison with the far–field dynamics. Thus, in the unscreened regime ($\kappa r_{ij}\ll 1$), the stream function of dipole $j$ is given by Eq. (14), leading to the Hamiltonian
$H_{\rm near}=\frac{1}{8\pi\eta_{s}}\sum_{i\neq j}\sigma_{i}\sigma_{j}\sin\!\big[2(\theta_{ij}-\alpha_{j})\big].$ (22)
For co–aligned dipoles ($\alpha_{j}=\alpha$), a global rotation sets $\alpha=0$, simplifying the angular structure. Differentiating Eq. (22) then yields explicit equations of motion that generate the $1/r$ near–field stresslet flows.
IV.2 Far–field Hamiltonian
In the screened regime ($\kappa r\gg 1$), the dipolar flow becomes vorticity free, as shown in Eq. (19). With no hydrodynamic torque acting on the dipole axes, orientations remain fixed and the dynamics reduce naturally to a position–only description. This irrotational structure, together with membrane incompressibility, makes the far field an intrinsically Hamiltonian regime: the pairwise interactions derive directly from the screened stream function. For well–separated dipoles ($\kappa r_{ij}\gg 1$), the screened stream function (18) yields
$H_{\rm far}=\frac{1}{2\pi\zeta_{\parallel}}\sum_{i\neq j}\frac{\sigma_{i}\sigma_{j}}{r_{ij}^{2}}\sin\!\big[2(\theta_{ij}-\alpha_{j})\big],$ (23)
The associated equations of motion,
$\dot{x}_{i}=\frac{1}{\sigma_{i}}\frac{\partial H}{\partial y_{i}},\qquad\dot{y}_{i}=-\frac{1}{\sigma_{i}}\frac{\partial H}{\partial x_{i}},$ (24)
generate the $1/r^{3}$ screened flows of Eq. (17). For co–aligned dipoles, both Hamiltonians reduce to simple functions of the relative polar angles, $H\propto\sum_{i\neq j}\sigma_{i}\sigma_{j}\sin(2\theta_{ij})/r_{ij}^{n}$ with $n=0$ or $2$.
V Collective dynamics of pusher and puller dipoles
We probe the dynamical consequences of the two Hamiltonians $H_{\rm far}$ and $H_{\rm near}$ by separately time-evolving ensembles of identical pushers and pullers ($\sigma_{i}=\pm 1$) with fixed, co–aligned orientations. In the far field, the flow is vorticity–free and orientations are naturally quenched, making $H_{\rm far}$ an exact position–based Hamiltonian. In the near field, where vorticity is finite, a Hamiltonian description applies only in the enforced quenched–orientation limit. Within these regimes, we contrast the collective organization produced by $H_{\rm far}$ and $H_{\rm near}$.
V.1 Simulation details
We integrate the $2N$ position variables $\{x_{i}(t),y_{i}(t)\}$ using a standard adaptive Runge–Kutta method, evolving Eq. (24) with either $H_{\rm near}$ or $H_{\rm far}$. Throughout this section we take $N=15$ and use identical random initial positions for both kernels: particles are sampled uniformly from a unit disc. Initial positions are indicated by open circles in Fig. 2, while filled black circles mark the final positions at $t=t_{\rm max}$. To quantify aggregation or dispersion, we monitor the mean pairwise separation
$\langle d_{ij}\rangle(t)=\frac{2}{N(N-1)}\sum_{i<j}\sqrt{\Delta x_{ij}^{2}+\Delta y_{ij}^{2}},$ (25)
where $\Delta x_{ij}=x_{i}-x_{j}$ and $\Delta y_{ij}=y_{i}-y_{j}$. A small softening parameter $\epsilon$ was added inside the squared distance to regularize close encounters without altering the large–scale flow. Energy conservation, $\dot{H}=0$ (between collisions) is verified to within a relative tolerance of $10^{-6}$, confirming numerical accuracy.
V.2 Near vs. far: aggregation and dispersion
The collective dynamics of active dipoles in an incompressible membrane depend crucially on whether interactions are mediated by the screened $H_{\rm far}$ or by the unscreened $H_{\rm{}_{n}ear}$. Figure 2 summarizes the behaviour for ensembles of $N=15$ pushers (top row) and pullers (bottom row).
Under the screened far–field Hamiltonian $H_{\rm far}$ [Figs. 2(a,b,e,f)], both motor types undergo a rapid collapse into compact aggregates: trajectories contract toward a single point and the mean pair separation $\langle d_{ij}\rangle(t)$ decreases sharply before saturating at a small value. This aggregation reflects the effectively attractive character of the $1/r^{3}$ screened dipolar flow, which reinforces head–to–tail alignment and produces a robust self–binding instability for both pushers and pullers.
In contrast, evolution under the unscreened near–field Hamiltonian $H_{\rm near}$ [Figs. 2(c,d,g,h)] exhibits a completely different trend. The $1/r$ stresslet decay is slower and we find that the ensemble does not collapse. Instead, particles spread into extended configurations and $\langle d_{ij}\rangle(t)$ grows monotonically. In this near–zone regime, the self–aggregation observed in the far field is vastly suppressed: although individual pairs can transiently attract or repel, they fail to form many–body stable compact clusters .
The similarity between the pusher and puller panels highlights a central result of this work: the nature of the hydrodynamic kernel, rather than the motor type alone, determines whether active dipoles aggregate or disperse. The far–field screened flow is irrotational and generates an attractive effective Hamiltonian, while the near–field unscreened flow disfavors collective collapse. For completeness, Fig. 3 shows a representative far–zone evolution with arbitrary orientations, illustrating that the same aggregation mechanism persists beyond the co–aligned limit. Representative far- and near-zone trajectories of co-aligned pusher and puller clusters are provided in the Supplementary Material (Figs. S1–S2). In all these examples, a soft harmonic repulsion is added to regularize close encounters and suppress particle overlap.
VI Summary and outlook
We have provided a concrete realization of a minimal hydrodynamic Hamiltonian framework for active force dipoles confined to an incompressible supported membrane. Starting from the Brinkman–regularized Stokes equation, we derived the screened incompressible Green’s tensor, identified its logarithmic near–field and algebraically screened far–field limits, and constructed the associated dipolar flows and stream functions. A key outcome of this analysis is the emergence of two distinct Hamiltonian constructions. In the far field, the screened flow is strictly vorticity–free, so dipole orientations remain naturally quenched; membrane incompressibility then yields an exact position–based Hamiltonian $H_{\rm far}$. In contrast, the near field possesses finite vorticity, precluding an orientation–dynamics–free description; a Hamiltonian formulation applies only in the enforced quenched–orientation limit, defining the effective Hamiltonian $H_{\rm near}$. Within these regimes, we contrast the collective organization generated by the two hydrodynamic Hamiltonians.
This comparison reveals a striking dichotomy. For identical pushers (and similarly for pullers), the far–field Hamiltonian $H_{\rm far}$ produces a strong self–binding instability, driving rapid collapse into compact clusters and a sharp decay of the mean pair separation. The unscreened near–field Hamiltonian $H_{\rm near}$ behaves qualitatively differently: it induces strong local rearrangements but no long–lived aggregates, and the ensemble disperses monotonically in time, rendering near–zone dynamics intrinsically non–aggregating. This contrast persists across motor types and initial orientations, showing that it is the structure of the hydrodynamic Hamiltonians—not merely the dipole sign—that controls collective organization. Moreover, our work establishes the important roles of membrane vorticity and incompressibility that leads to the emergence of such Hamiltonian descriptions. These results offer a framework for interpreting more complex membrane rheologies. Allowing membrane compressibility opens a longitudinal mobility channel that breaks the Hamiltonian structure and can enhance aggregation for pullers [24] while producing mixed responses for pushers. Introducing odd (Hall) viscosity adds antisymmetric stresses that drive chiral transport and circulating cluster states. An important open question is the extent to which such effects can be incorporated within generalized Hamiltonian frameworks, [24]. Coupling the membrane to curvature, shape fluctuations, or biochemical activity further broadens the phenomenology, enabling a range of richer and biologically relevant collective behaviors, some of which may admit generalized Hamiltonian descriptions based on symmetry considerations. More broadly, incompressible supported membranes provide a minimal yet dynamically rich example in which explicit Hamiltonian descriptions of active particles emerge under suitable conditions. This perspective motivates experimental tests in reconstituted active films, where screening lengths and membrane viscosities can be precisely tuned, as well as theoretical extensions of Hamiltonian and symmetry-based approaches to active matter at interfaces.
Acknowledgements.
We are very thankful to Sarthak Bagaria, Naomi Oppenheimer, Haim Diamant, Michael D. Graham, and Mark Henle. S.K. acknowledges support from an Institute fellowship at Birla Institute of Technology and Science, Pilani (Hyderabad Campus). R.S. is supported by a DST INSPIRE Faculty fellowship, India (Grant No. IFA19-PH231), NFSG, and an OPERA Research Grant from Birla Institute of Technology and Science, Pilani (Hyderabad Campus).
Data Availability
The numerical implementations used in this study are available from the authors upon reasonable request.
Appendix A Supplementary Figures
(a) Pushers (far zone)
(b) Pullers (far zone)
(c) Pushers (near zone)
(d) Pullers (near zone)
Figure 4: Far- and near-zone dynamics of co-aligned clusters of twelve active particles in an incompressible membrane. Columns correspond to pushers (left) and pullers (right). Particles are initialized at random locations within a disc. Initial positions are shown in green (pushers) and yellow (pullers), final positions in red and orange, respectively, with trajectories shown in grey. A soft harmonic repulsion is included to prevent particle overlap.